OpenStudy (anonymous):

Find the radius of the circle with these three points: (-4,9), (2,9), and (-5,8) and center: (-1,5).

OpenStudy (anonymous):

The center tells you alot: That gives you this without any work: \[(x+1)^2+(y-5)^2=r^2\]. From here you can plug in x and y and solve directly for r.

OpenStudy (anonymous):

which x and y

OpenStudy (anonymous):

Any of them :) They all have to satisfy the equation. So they should all give the same r.

OpenStudy (anonymous):

Do you have to foil both sides?

OpenStudy (anonymous):

You don't have to. For example try the point (-4,9). You get: \[((-4)+1)^2+(9-5)^2=r^2\] \[(-3)^2+(4)^2=r^2\] \[9+16=r^2\] \[25=r^2\] r=5

OpenStudy (anonymous):

would it be -5

OpenStudy (anonymous):

The radius is always defined as a positive number. But 5 is correct :)

OpenStudy (anonymous):

thanks for the help malevolence19 {;

OpenStudy (anonymous):

Do you know how to solve Find the center of the circle defined by the equation 6x^2+6y^2+48x-36y-144=0

OpenStudy (anonymous):

Yes. You complete the square. \[6(x^2+8x)+6(y^2-6y)=144\]. (rewriting) Then: \[6(x^2+8x+16-16)+6(y^2-6y+9-9)=144\] Then factoring: \[6(x+4)^2+6(y-3)^2-16(6)-9(6)=144\] \[6[(x+8)^2+(y-3)^2]=204\] \[(x+8)^2+(y-3)^2=34\]

OpenStudy (anonymous):

Err those last lines should say (x+4)

OpenStudy (anonymous):

Wrong. So 144+54+96=294. Then dividing by 3 gives 49. So you should have: (x+4)^2+(y-3)^2=49. Sorry I just failed so hard >.<